# Machine learning for hydrodynamic stability

### v2

@@ 3. Validation of the classification strategy
The classification boundary in Figure 7 is consistent with estimates of the critical Reynolds number $\mathcal{R}_{*}$ for unperturbed flow ($\delta=0$) that have been published in the literature. Results for three studies reported in Table II in [7] suggest that $\mathcal{R}_{*}$ is between 216 and 219, but the authors also remark that “the exact value of the critical Reynolds number is difficult to fix because it depends on grid resolution, but it is close to 216.” This remark is supported by the eigenvalue results in [10, Fig. 4] showing that computed estimates of the critical eigenvalue governing linear stability move to the left as the spatial resolution is increased. To assess the sensitivity to discretisation parameters and validate our surrogate model, a second set of 50 classification data points were randomly generated close to the classification boundary with a refined spatial grid (level 6 in ifiss). In these calculations, the dimension ${n_{x}}$ of the algebraic system is increased from 23,842 to 94,146 and ${n_{t}}$, the number of time steps, is increased from 1200 to 1600. The results are shown in Figure 8. The only two parameter combinations that are misclassified by the coarse grid surrogate model have a perturbation of magnitude close to $10^{-4}$. Looking carefully at the results in the left plot in Figure 7 one can see a lot of uncertainty in the classification of flows with perturbation of this magnitude. We did not investigate this further. We note however that if the magnitude of the perturbation is increased to $5\times 10^{-3}$ then the classification results in Figure 7 suggest that the flow is not symmetric for all Reynolds numbers in the range considered.
@@ 4. Classification of instabilities in thermal convection
The broad applicability of our classification strategy will be demonstrated by exploring the bifurcation structure of two representative problems that arise when modelling buoyancy driven flow. The first case we consider is a symmetry-breaking pitchfork bifurcation and the second is a Hopf bifurcation from a steady flow solution to a periodic flow. We note that both problems are members of the set of canonical bifurcation test problems discussed by Wubs & Dijkstra in [23, Chap. 10].
@@ 4.1. A Rayleigh–Bénard convection problem
Following their discovery by Bénard in the laboratory in 1900, the development of recirculating convection cells has been extensively studied: both theoretically, see [5, Chap. 6], and computationally, see Gelfgat [12]. A state-of the-art algorithm that combines Newton’s method with a simple deflation strategy can be found in Boulle et al. [2].
@@
$k(t)={1\over 2}\int_{\Omega}\vec{u}(\cdot,t)\cdot\vec{u}(\cdot,t)\,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\rm{d}\Omega},$ (9)
@@
[…]
Figure 10. Step size history (left) and kinetic energy (right) for ${Ra}=1530$, $\delta=0$. The vertical lines refer to specific time steps that are highlighted in the text.
@@
  • The first phase is the development of the equilibrium solution as the horizontal walls are heated/cooled from zero to their final values of ±0.5plus-or-minus0.5$\pm$ 0.5± 0.5.555The visible jumps in the time step after 30 and after 270 steps correspond to an TR–AB2 averaging step. If averaging is not done then the time step stagnates before the equilibrium solution is established. Averaging was done every 30 steps until step 300, but was not invoked for the final 500 steps. A tiny jump in the kinetic energy between time steps 33 and 66 (too small to feature on the plot) is associated with the motion of the fluid needed to initiate the transition to the equilibrium state. A linear temperature profile is clearly established after 20 time units and 210 time steps.
@@
The dataset $SS\SS$ was generated by taking 60 uniformly spaced Rayleigh numbers between 1300 and 1600 and 7 logarithmically spaced perturbations between $10^{-16}$ and $10^{-2}$. A flow problem is classified as “bifurcated” $(\ell^{j}=1)$ if the final kinetic energy was greater than ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2\times 10^{-7}}$ (this is clearly the case for the parameter combination in Figure 10). With this test in place, 254 of the data points $SS\SS$ were labelled as a system in the conducting state and 166 were classed as asymmetric. The precise classification is shown in Figure 11. Sampling the trained neural network at 200,000 points leads to the prediction of the bifurcation boundary shown in the right plot. From this picture one would predict that the critical Rayleigh number for perfect uniform heating of the cavity bottom is in the interval $(1510,1520)$. The results also suggest that the critical Rayleigh number will be less than 1500 if the applied heating is skewed by a measurable amount $\geq 10^{-3}$, one way or the other. The results also indicate that if the perturbation is larger than $\geq 10^{-2}$, then the critical Rayleigh number will be less than 1300.
@@ 4.2. A differentially heated cavity problem
The second problem is motivated by laboratory experiments originally performed by Elder [8]. The experiments were performed over 50 years ago by observing the motion of aluminium powder suspended in viscous oil with a Prandtl number 1000. The aim here is to extend our understanding with a view to assessing the stability of the laboratory configuration for a variety of fluids with Pr ranging from 7.1 (water) to 1000 (glycerol). We note that the bifurcation structure for fluids with low Prandtl numbers has also been extensively studied over the past 25 years, see Gelfgat et al. [13]. Elder’s original experimental set up can be modelled as a (tall) 0.051:1 aspect ratio cavity with insulated horizontal boundaries $y=0$ and $y=1$. The left boundary $x=0$ is hot and the right boundary $x=0.051$ is cold. If the imposed temperature difference is sufficiently small then a steady flow solution will be established with a clockwise primary flow recirculation.
@@
If the Rayleigh number is above a critical value then the steady solution loses stability via a Hopf bifurcation. The associated Rayleigh number that was identified from the laboratory experiments is $2.26\times 10^{9}\pm 30\%$ [^7] The definition of the Rayleigh number in [8] is scaled by a factor of ${d}^{3}$., the large uncertainty is attributed to the difficulty of detecting the onset of a very weak secondary flow. A precise critical value of $2.88\times 10^{9}$ is identified in [23, Sect. 10.4] by carefully tracking eigenvalues. Computational results for a Rayleigh number just above this value typically evolve to a periodic flow solution with small “ripples” along the vertical centreline $y=0.5$. Such an instability is apparent in the volume averaged kinetic energy history for a Rayleigh number of $2.89\times 10^{9}$ that is presented in Figure 12, wherein a stable oscillation with a period of 683 time units can be clearly identified after the first 1000 steps of integration. In contrast, when the computed kinetic energy is plotted on the same scale for a slightly smaller Rayleigh number, say Ra = $2.87\times 10^{9}$, there are no oscillations visible after 1000 time steps, see Figure 13.
Figure 12. Evolution of the kinetic energy for a differentially heated cavity with Ra=$2.89\times 10^{9}$ and Pr=1000: time steps 120–2000 (left), time steps 1200–2000 (right).
Figure 13. Evolution of the kinetic energy for a differentially heated cavity with Ra=$2.87\times 10^{9}$ and Pr=1000: time steps 120–2000 (left), time steps 1200–2000 (right).
@@
  • The first phase is the development of the primary recirculation as the vertical walls are heated/cooled from zero to their final values of ±$\pm 0.5$. There is no analytical solution in this case—even a small temperature difference will generate a nontrivial flow in the cavity. The primary recirculation is fully developed after 100 time steps, which translates to about 200 units of time.
@@
The parameters defining the bifurcation problem that is the focus here are the Rayleigh number (Ra) and the Prandtl number (Pr). The dataset $SS\SS$ was generated for 7 specific values of Pr; namely, 7.1, 71, 140, 195, 250, 500 and 1000. We computed results for different values of Ra  between $2.81\times 10^{9}$ and $3.09\times 10^{9}$. A flow problem is classified as “bifurcated” $(\ell^{j}=1)$ if, when looking at the final 800 time steps of the computation, the ratio of the difference between the maximum and the minimum energy $k$ divided by the mean value of $k$ is greater than ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}6\times 10^{-4}}$ (as is the case for the parameter combination in Figure 12). With this test in place, 45 of the data points $SS\SS$ were labelled as a system that reached a steady state and 49 points were classed as being oscillatory. The precise classification is shown in Figure 14.
@@
  • Both the raw data for Pr=1000 and the bifurcation point predicted by the trained neural network are in close agreement with the results in [23].
@@
Confirmation of these observations clearly requires an adaptive strategy. One needs to refine the surrogate approximation by generating additional data points in regions of parameter space to reduce the interpolation error where the bifurcation boundary is most uncertain. The design of a semi-rigorous adaptive strategy—where the interpolation error is balanced by the PDE approximation error—is left as future work.
5. Concluding remarks
@@
While reliably trained neural networks offer the possibility of transformative technology in particular areas of science, a convincing case for replacing physics-based modelling by contemporary machine-learning strategies is still to be established at this time. The strategy presented herein has a very different flavour: the interpolating potential of a trained neural network is utilised in a role where it is known to be very effective.
@@
Our use of a robust adaptive time-stepping strategy to follow the evolution of a flow from a quiescent state to a final steady or periodic state is a key feature of our strategy. Our impression is that integration of the Navier–Stokes equations in time has invariably been done by researchers in the past using fixed time steps. While fixed time-stepping methods can be effective in modelling the growth or otherwise of perturbations of the equilibrium flow, it cannot be used to model a laboratory experiment from start to finish.
@@
While we have concentrated on tracking the first bifurcation point in this work, there is no reason why the methodology cannot be to adapted to locate a succession of branch points. All the flow problems considered in this work are representative of problems where a classical linearised stability analysis is predictive of reality in the sense of being able to reproduce results observed in laboratory experiments. One interesting future direction would be to consider applying the methodology to flow problems where transient growth of instabilities limits the applicability of classical eigenvalue analysis.

### v4

@@ 3.2. Blind data test results
An important observation is that the probability distribution generated by the blind dataset surrogate is perfectly consistent with estimates of the critical Reynolds number $\mathcal{R}_{*}$ for unperturbed flow ($\delta=0$) that have been published in the literature. Results for three studies reported in Table II in [Drikakis(1997)] suggest that $\mathcal{R}_{*}$ is between 216 and 219, but the authors also remark that “the exact value of the critical Reynolds number is difficult to fix because it depends on grid resolution, but it is close to 216.” This remark is supported by the eigenvalue results in [Elman and Silvester(2018), Fig. 4] showing that computed estimates of the critical eigenvalue governing linear stability move to the left as the spatial resolution is increased.
3.3. Validation of the blind test results
[…]
Figure 9. Symmetric step channel flow showing surrogate predictions from trained neural network using 1476 uniformly spaced data points. Colour variation highlights the bifurcation boundary solution (left plot). Estimated probability of observing a bifurcated solution for a fixed perturbation (right plot).
Comparing the results in Figures 8 and 9 we note four distinctive features:
  • The zone of uncertainty in the predicted bifurcation boundary with $\delta<10^{-3.5}$ is significantly reduced in the fine grid surrogate (no surprise, there are more sample points in the uncertain region).
  • The fine grid surrogate introduced more uncertainty in the position of the boundary for perturbations $\delta$ larger than $10^{-3.5}$.
  • The coarse grid surrogate gives an excellent prediction of the probability of observing a bifurcation for the two parameter values that are highlighted.
  • With a visually imperceptible inlet profile perturbation of magnitude of $10^{-3.5}$ the critical Reynolds number decreases from the unperturbed value of 220 to a value close to 210. This is a striking and surprising result.
@@ 4. Classification of instabilities in thermal convection
The broad applicability of our classification strategy will be demonstrated by exploring the bifurcation structure of two representative problems that arise when modelling buoyancy driven flow. The first case we consider is a symmetry-breaking pitchfork bifurcation and the second is a Hopf bifurcation from a steady flow solution to a periodic flow. We note that both problems are members of the set of canonical bifurcation test problems discussed by Wubs & Dijkstra in [Wubs and Dijkstra(2023), Chap. 10]. The PDE system that we use to model such flows is given by
${\partial{\vec{u}}\over\partial t}+{\vec{u}}\cdot\nabla{\vec{u}}-{\nu}\,\nabla^{2}{\vec{u}}+\nabla p$ $=\theta{\vec{j}},$ (10)
$\nabla\cdot{\vec{u}}$ $=0,$
${\partial\theta\over\partial t}+{\vec{u}}\cdot\nabla\theta-{\epsilon}\,\nabla^{2}\theta$ $=0,$
where $\theta$ is the temperature, and we observe that the Navier–Stokes equations (6) are being driven by the buoyancy forcing term $\theta{\vec{j}}$ acting downwards. The fluid velocity $\vec{u}$} is always set to zero on the domain boundary. Working in dimensionless units the diffusion parameters $\nu$ and $\epsilon$ in (10) can be expressed in terms of the characteristic Rayleigh number Ra and Prandtl number Pr via
$\nu=\sqrt{{\hbox{Pr}\over{Ra}}},$ $\quad\epsilon={1\over\sqrt{\hbox{Pr}{Ra}}}.$ (11)
Note that the two diffusion parameters are close to each other when Pr is close to 1; for example, when modelling the thermal convection of air, but are significantly different for fluids with Pr $\gg 1$.
@@ 4.1. A Rayleigh–Bénard convection problem
Following their discovery by Bénard in the laboratory in 1900, the development of recirculating convection cells has been extensively studied: both theoretically, see [Drazin(2002), Chap. 6], and computationally, see Gelfgat [Gelfgat(1999)]. A state-of the-art algorithm that combines Newton’s method with a simple deflation strategy can be found in Boulle et al. [Boullé et al.(2022)].
@@
$k(t)={1\over 2}\int_{\Omega}\vec{u}(\cdot,t)\cdot\vec{u}(\cdot,t)\,\rm{d}\Omega,$ (12)
@@
$\begin{aligned} T(\vec{x},t)&=\left\{{1\over 2}+\delta\cdot\sin(2\pi x)\right\}(1-e^{-10t}),&0\leq x\leq 1,y=0,t>0,\\ T(\vec{x},t)&=-{1\over 2}(1-e^{-10t}),&0\leq x\leq 1,y=2,t>0.\end{aligned}$ (13)
Numerical solutions are computed with the default ifiss spatial approximation $Q_{2}$$Q_{1}$$Q_{2}$ of the Boussinesq flow equations (biquadratic temperature) [Elman et al.(2014b), Sect. 11.3]. Labelled data is generated by discretising the cavity domain into a uniform grid of $32\times 64$ square elements. The dimension $n_{x}$} of the resulting algebraic system in this case is 27,300. The linearised TR–AB2 integrator described in [Elman et al.(2014b), Sect. 11.2] (one ${n_{x}}\times{n_{x}}$ sparse matrix solve per time step) is used for generating the numerical solution. The linear equation systems are near singular (in exact arithmetic there is a constant pressure vector in the null space). Numerically singular systems that arise in the course of the time integration are dealt with by constructing and solving an augmented (regularised) system, see Bochev & Lehoucq [Bochev and Lehoucq(2005)]. The initial time step is 1e-9. The stabilised TR–AB2 integrator is run for a fixed number $n_{t}$} steps with the time accuracy tolerance parameter set to 1e-6. When generating labelled data, $n_{t}$} is set to 800.
Figure 11. Step size history (left) and kinetic energy (right) for Ra=1530{Ra}=1530, $\delta=0$. The vertical lines refer to specific time steps that are highlighted in the text.
@@
  • The first phase is the development of the equilibrium solution as the horizontal walls are heated/cooled from zero to their final values of $\pm 0.5$. [^5] The visible jumps in the time step after 30 and after 270 steps correspond to an TR–AB2 averaging step. If averaging is not done then the time step stagnates before the equilibrium solution is established. Averaging was done every 30 steps until step 300, but was not invoked for the final 500 steps. A tiny jump in the kinetic energy between time steps 33 and 66 (too small to feature on the plot) is associated with the motion of the fluid needed to initiate the transition to the equilibrium state. A linear temperature profile is clearly established after 20 time units and 210 time steps.
@@
Figure 13. Rayleigh–Bénard convection showing surrogate predictions from trained neural network using 420 data points. Colour variation highlights the bifurcation boundary solution (left plot). Estimated probability of observing a bifurcated solution for a fixed perturbation (right plot).
The dataset $S\SS$ was generated by taking 60 uniformly spaced Rayleigh numbers between 1300 and 1600 and 7 logarithmically spaced perturbations between $10^{-16}$ and $10^{-2}$. A flow problem is classified as “bifurcated” $(\ell^{j}=1)$ if the final kinetic energy was greater than $2\times 10^{-7}$ (this is clearly the case for the parameter combination in Figure 11). With this test in place, 254 of the data points $S\SS$ were labelled as a system in the conducting state and 166 were classed as asymmetric. The precise classification is shown in Figure 12. Sampling the trained neural network at 200,000 points leads to the prediction of the bifurcation boundary shown in the right plot.
The visualisation of the predicted bifurcation boundary in Figure 13 follows the strategy described in section 3.2. As in that case, the uniformly sharp transition is the prominent feature. From this set of labelled data, one can be confident that the critical Rayleigh number for perfect uniform heating of the cavity bottom is in the interval $(1510,1520)$. In this case the visually imperceptible temperature profile perturbation of magnitude of $10^{-3.5}$ decreases the critical Rayleigh number from the unperturbed value to a value close to 1505. Comparing the fixed perturbation probability profiles with those in Figure 9, we see that the effect of the perturbation is similar in both cases. This is probably to be expected.
@@ 4.2. A differentially heated cavity problem
The second problem is motivated by laboratory experiments originally performed by Elder [Elder(1965)]. The experiments were performed over 50 years ago by observing the motion of aluminium powder suspended in viscous oil with a Prandtl number 1000. The aim here is to extend our understanding with a view to assessing the stability of the laboratory configuration for a variety of fluids with Pr ranging from 7.1 (water) to 1000 (glycerol). We note that the bifurcation structure for fluids with low Prandtl numbers has also been extensively studied over the past 25 years, see Gelfgat et al. [Gelfgat et al.(1999)]. Elder’s original experimental set up can be modelled as a (tall) 0.051:1 aspect ratio cavity with insulated horizontal boundaries $y=0$ and $y=1$. The left boundary $x=0$ is hot and the right boundary $x=0.051$ is cold. If the imposed temperature difference is sufficiently small then a steady flow solution will be established with a clockwise primary flow recirculation.
Numerical solutions are computed with $Q_{2}$$Q_{1}$$Q_{2}$ spatial approximation using ifiss. Labelled data is generated by discretising the cavity domain into a nonuniform grid of $32\times 256$ rectangular elements that is geometrically stretched away from the walls. [^6] The grid is defined by setting the horizontal stretch factor to 2 and the vertical stretch factor to 1.5. The dimension $n_{x}$} of the resulting algebraic system in this case is 108,516. The TR–AB2 integrator is run in fully nonlinear mode with a second-order linearisation step followed by two fixed-point iterations as in section 3. The integrator is run for a fixed number $n_{t}$} steps with the time-accuracy tolerance parameter set to 2e-5. When generating labelled data $n_{t}$} is set to 2000. To model the action of heating the walls in a laboratory setting the temperature variation on the vertical walls is ramped up in time, multiplying the steady-state constant boundary temperature profile by the lifting function $1-e^{-10t}$.
If the Rayleigh number is above a critical value then the steady solution loses stability via a Hopf bifurcation. The associated Rayleigh number that was identified from the laboratory experiments is $2.26\times 10^{9}\pm 30\%$ [^7] The definition of the Rayleigh number in [Elder(1965)] is scaled by a factor of ${d}^{3}$., the large uncertainty is attributed to the difficulty of detecting the onset of a very weak secondary flow. A precise critical value of $2.88\times 10^{9}$ is identified in [Wubs and Dijkstra(2023), Sect. 10.4] by carefully tracking eigenvalues. Computational results for a Rayleigh number just above this value typically evolve to a periodic flow solution with small “ripples” along the vertical centreline $y=0.5$. Such an instability is apparent in the volume averaged kinetic energy history for a Rayleigh number of $2.89\times 10^{9}$ that is presented in Figure 14, wherein a stable oscillation with a period of 683 time units can be clearly identified after the first 1000 steps of integration. In contrast, when the computed kinetic energy is plotted on the same scale for a slightly smaller Rayleigh number, say Ra = $2.87\times 10^{9}$, there are no oscillations visible after 1000 time steps, see Figure 15.
@@
[…]
@@
  • The first phase is the development of the primary recirculation as the vertical walls are heated/cooled from zero to their final values of $\pm 0.5$. There is no analytical solution in this case—even a small temperature difference will generate a nontrivial flow in the cavity. The primary recirculation is fully developed after 100 time steps, which translates to about 200 units of time.
@@
The parameters defining the bifurcation problem that is the focus here are the Rayleigh number (Ra) and the Prandtl number (Pr). The dataset $S\SS$ was generated for 7 specific values of Pr; namely, 7.1, 71, 140, 195, 250, 500 and 1000. We computed results for different values of Ra between $2.81\times 10^{9}$ and $3.09\times 10^{9}$. A flow problem is classified as “bifurcated” $(\ell^{j}=1)$ if, when looking at the final 800 time steps of the computation, the ratio of the difference between the maximum and the minimum energy $k$ divided by the mean value of $k$ is greater than $6\times 10^{-4}$ (as is the case for the parameter combination in Figure 14). With this test in place, 45 of the data points $S\SS$ were labelled as a system that reached a steady state and 49 points were classed as being oscillatory. The precise classification is shown in Figure 16.
@@
  • Both the raw data for Pr=1000 and the bifurcation point predicted by the trained neural network are in close agreement with the results in [Wubs and Dijkstra(2023)].
@@
Confirmation of these observations can be realised by using an adaptive strategy. The idea is to refine the surrogate approximation by generating additional data points in regions of parameter space to reduce the interpolation error where the bifurcation boundary is most uncertain. One such approach is developed in [Singh and Silvester(2025)], where the classifier output provides a probabilistic measure of flow stability, and the Shannon entropy of these predictions is employed as an uncertainty indicator. A flow-based deep generative machine learning strategy can then be trained to approximate a probability density function that concentrates sampling in regions of high entropy, thereby directing computational effort toward the evolving bifurcation boundary. This coupling between classification and generative modelling establishes a feedback-driven adaptive learning process analogous to error indicator based refinement in contemporary PDE solution strategies.
5. Concluding remarks and critical assessment
@@
While reliably trained neural networks offer the possibility of transformative technology in particular areas of science, a convincing case for replacing physics-based modelling by contemporary machine-learning strategies is still to be established at this time. The strategy presented herein has a very different flavour: the interpolating potential of a trained neural network is utilised in a role where it is known to be state-of-the-art. The bifurcation tracking strategy is not perfect however. Here is a list of factors that need to be considered before adopting it:
  • Reproducibility is an obvious issue in machine learning. The loss function is rarely reduced to zero, even after intensive training. This means that “converged” learned weights in the surrogate model will be different if the training is repeated with a different realisation of the random initial weights. While this is a significant problem in a deep learning setting, it is less of an issue when training a shallow neural network with a relatively small number of parameters.
@@
  • The physics needs to be faithfully represented by the discretised mathematical model if the results are to be predictive in the sense of being able to reproduce results observed in practical experiments. Spatial approximation needs to be done carefully with either mesh refinement checks to confirm accuracy or refinement determined by reliable a posteriori error control. Our use of a robust adaptive time-stepping strategy to follow the evolution of a flow from a quiescent state to a final steady or periodic state is a notable feature of our strategy. Integration of the Navier–Stokes equations in time has typically been done in the past using fixed time steps. While fixed time-stepping methods can be effective in modelling the growth or otherwise of perturbations of the equilibrium flow, they cannot be used to model a laboratory flow experiment from start to finish.
  • The identification of a sensible bifurcation detection strategy is problem dependent and requires physical insight. For the three cases in this study, prior testing of alternatives was done to identify a suitable measure of departure from the equilibrium state and inform the choice of tolerance for labelling the training data. The detection process needs to be robust in the sense that the distinctive features of the bifurcation picture are independent of the precise choice of tolerance. This can be easily checked—one simply needs to rerun the training process. The expensive step of generating the raw data does not need to be repeated.
@@
While we have limited our attention to two-dimensional bifurcation problems in this work, our machine learning strategy generalises effectively to high-dimensional parametric problems. Moreover, while we have concentrated on tracking the first bifurcation point, there is no reason why the methodology cannot be to adapted to locate a succession of branch points. All the flow problems considered in this work are representative of problems where a classical linearised stability analysis is predictive of reality in the sense of being able to reproduce results observed in laboratory experiments. One interesting future direction would be to consider applying the methodology to flow problems where transient growth of instabilities limits the applicability of classical eigenvalue analysis. Another interesting extension of this work would be to incorporate labels generated from laboratory experiments into the training process.